Show that the closed sphere with centre (2,3,7) and radius 10 in R^3 is contained in the
open cube P = {(x, y, z) :! x − 2 !<11, !y − 3! <11, !z − 7! <11}..
1
Expert's answer
2020-03-17T15:45:53-0400

We can parametrize a sphere as the set of all points "(x,y,z)" such that "\sqrt{(x-2)^2+(y-3)^2+(z-7)^2}\leq 10"

"\\
\implies (x-2)^2+(y-3)^2+(z-7)^2\leq 100"

As each term here is a square, they are nonnegative, so we get "3" simultaneous inequalities

"(x-2)^2\leq 100<121\\
(y-3)^2\leq 100<121\\
(z-7)^2\leq 100<121\\
\implies |x-2|<11\ \&\ |y-3|<11\ \&\ |z-7|<11\"

Thus, any "(x,y,z)" belonging to the closed sphere belongs to "P", the open cube.

Thus, the closed sphere is contained in the open cube.